# VTechAGP: An Academic-to-General-Audience Text Paraphrase Dataset and Benchmark Models

### v1

@@ 6 Conclusion
We created VTechAGP, an academic-to-general-audience document-level translation benchmark dataset, which consists of 4,938 academic and general audience abstract pairs with their corresponding auxiliary information such as title, subject terms, etc. We explore several SOTA LLMs to establish the baseline performance on the text paraphrasing task. We propose a new model, DSPT5, which includes dynamic soft prompt generation with hybrid loss during the training phase and a new crowd-sampling strategy in the inference stage. We evaluate the datasets and models from the perspective of document-level embedding-based, word-based, end-to-end metrics, simplicity, diversity, readability, toxicity and human evaluation. Extensive experimental results on 8 different colleges of VTechAGP show that DSPT5 achieves comparable results with other SOTA benchmark models.
@@ 7 Limitations
Based on the exploration of the datasets in Sec. B and the experiments of the benchmark models in Sec. 5, we identify the following limitations (remaining challenges) and future directions:
(1) Dataset: Although VTechAGP is a recent and distinctive dataset, it suffers in terms of size, with a total of 4,938 document pairs. This is due to the fact that the policy of including a general audience version of the abstract for all theses and dissertations submitted to VTechWorks began to be implemented in 2016. However, this aspect of the limitation should be gradually overcome as VTechWorks continues to collect and maintain this collection of documents. With the increase in graduate student enrollment at Virginia Tech in recent years, it is safe to assume that VTechAGP will continue to grow in the coming years. In addition, VTechAGP consists of pairs of abstracts carefully written by students, encapsulating their work during their time in graduate school. With about 50% of the corpus composed of Ph.D. dissertations and the other half composed of M.S. theses, we believe that VTechAGP should achieve good quality due to the amount of time and effort put into the creation process. As the size of the dataset grows, VTechAGP will become more robust for text paraphrasing tasks in the near future.
(2) Model: While transformer-based LLMs retain their popularity for text generation tasks, our experiment results in Sec. 5 show that there is still a huge room for further improvement of LLMs in academic to general audience text paraphrasing tasks. With the rapid development of LLMs, more robust and efficient LLMs are emerging. For example, the recently released model Gemini Team et al. (2023) shows great potential and competitive results compared to ChatGPT and LLaMA2 in natural language understanding-related tasks. In addition, some recently released lightweight decoder-only models (e.g., Phi-2 Li et al. (2023)) also exhibit great potential for text generation tasks. As the VTechAGP dataset grows, it will be able to support fine-tuning for more powerful open-source LLMs (i.e. Phi-2, Gemini).
@@ Appendix A Dataset Comparison
Breath-first search (BFS) is a fundamental building block in many graph-based applications, but it is difficult to optimize for a field-programmable gate array (FPGA) due to its irregular memory-access patterns. Prior work, based on hardware description languages (HDLs) and high-level synthesis (HLS), address the memory-access bottleneck of BFS by using techniques such as data alignment and compute-unit replication on FPGAs. The efficacy of such optimizations depends on factors such as the sparsity of target graph datasets. Optimizations intended for sparse graphs may not work as effectively for dense graphs on an FPGA and vice versa. This thesis presents two sets of FPGA optimization strategies for BFS, one for near-hypersparse graphs and the other designed for sparse to moderately dense graphs. For near-hypersparse graphs, a queue-based kernel with maximal use of local memory on FPGA is implemented. For denser graphs, an array-based kernel with compute-unit replication is implemented. Across a diverse collection of graphs, our OpenCL optimization strategies for near-hypersparse graphs delivers a 5.7x to 22.3x speedup over a state-of-the-art OpenCL implementation, when evaluated on an Intel Stratix 10 FPGA. The optimization strategies for sparse to moderately dense graphs deliver 1.1x to 2.3x speedup over a state-of-the-art OpenCL implementation on the same FPGA. Finally, this work uses graph metrics such as average degree and Gini coefficient to observe the impact of graph properties on the performance of the proposed optimization strategies.
@@ Appendix B Dataset Analysis
VTechAGP contains a total of 4,938 document pairs (academic abstract and general audience abstract), where 52.8% of the documents are from Ph.D. dissertations and 47.2% of the documents are from Master’s theses. Table 1 shows the data statistics of VTechAGP across eight different colleges. VTechAGP is divided into eight colleges because data from different colleges exhibits distinct statistics. In addition, the dataset divided into colleges can also support potential cross-domain research tasks and provide the convenience for researchers focusing on a specific domain in the future. The document pairs are categorized based on colleges since authors from the same college may share some common terminology and background knowledge. Therefore, these categories should exhibit similar statistics. Table 1 reports some basic dataset statistics, including the number of documents for each college, the average number of sentences for each document, and the average sentence length for the source and target texts, respectively. We also report textual and lexical diversity (MTLD), which reflects the average number of words in a row for which a certain type-token ratio is maintained McCarthy (2005). To better illustrate the difference in readability between the source and target texts, we report Readability Consensus, [^6] https://pypi.org/project/textstat/ which is a combined evaluation of Flesch Kincaid Grade, Flesch Reading Ease, SMOG Index, Coleman-Liau Index, Automated Readability Index, Dale-Chall Readability Score, Linsear Write Formula, and Gunning FOG Formula. From Table 1, we observe that the general audience abstracts have fewer sentences and shorter sentence lengths compared to the academic abstracts. Furthermore, academic abstracts have better lexical diversity than general audience abstracts, except for the College of Architecture, Arts and Design, and the College of Business. In general, academic abstracts have a higher level of comprehension difficulty compared to general audience abstracts, except the College of Engineering. Analysis for the College of Engineering, which represents 49.6% of the total dataset, is discussed below.
@@
Department # Doc. Readability Consensus
Aerospace Eng. 181 14th-15th / 14th-15th
Biomedical Eng. 234 13th-14th / 10th-11th
Building Construction 19 15th-16th / 15th-16th
Chemical Eng. 46 16th-17th / 11th-12th
Civil Eng. 234 16th-17th / 15th-16th
Computer Sci. 392 14th-15th / 15th-16th
Electrical & Computer Eng. 749 15th-16th / 15th-16th
Eng. Education 31 17th-18th / 16th-17th
Environmental Sci. & Eng. 23 15th-16th / 14th-15th
Industrial & Systems Eng. 86 16th-17th / 15th-16th
Materials Sci. & Eng. 72 16th-17th / 11th-12th
Mechanical Eng. 349 15th-16th / 14th-15th
Mining Eng. 35 15th-16th / 15th-16th
The distinct category (College of Engineering) contains the largest number of document pairs, representing 49.6% of the total dataset, due to the large number of students from 13 different departments in the College of Engineering. To better understand the characteristics of this category, we divided the documents in the College of Engineering based on their corresponding departments. Details are given in Table 8. Thus, we expect document pairs within this college to be more diverse compared to others. From Table 8, we observe that documents from most departments in the College of Engineering have a similar readability consensus between academic abstracts and general audience abstracts. The only outlier is the Department of Computer Science, which shows that general audience abstracts are more difficult to understand than academic abstracts. On the other hand, the Department of Electrical and Computer Engineering alone represents 15.2% of the total dataset, while maintaining an above-average readability consensus for its general audience abstracts, thus contributing to the higher readability consensus for the College of Engineering. We conjecture as follows: 1) Various complex terminologies from some research fields are inevitably retained in general audience abstracts because they cannot be replaced by simple terms. 2) Concepts in certain areas of engineering research are difficult to explain or rewrite into short and concise sentences. 3) Some engineering students are not as good at simplification writing as students from other colleges.
@@ Appendix C Parameter Settings
For the hyperparameters and configuration of DSPT5, we implement DSPT5 in PyTorch and optimize it with the AdamW optimizer. For pre-trained LLMs, we retrieve the generated output through their API by giving the prompt "Generate another version of the provided document for the general audience". We use grid search to tune the hyperparameters. The learning rate is $\in\left\{5e-2,5e-3,5e-4,5e-5\right\}$. The contrastive loss weight $\lambda$ in Equ. 8 and the alignment function weight $\gamma$ in Equ. 10 are chosen from 0.1 to 0.9 with a step size of 0.2. The number of candidates $C_{n}$ in Sec. 4.3 is chosen from $\left\{4,8,16\right\}$. The batch size is 4. The length of the source text, fixed prompt template, and keywords are 512, 16, and 16, respectively. The fixed prompt template used for baselines is: ‘Generate another version of the provided document for general audiences.’. We choose the hyperparameters based on the validation set. Because the dataset is not a balanced dataset, where College of Engineering takes up most of the datasets. We split a validation set in the domain of College of Engineering documents. This is because as shown in Table 1, it is not realistic to have a validation set in the Business domain, which has only 63 documents in total. The size of the dataset will only grow as new master’s theses or PhD dissertations are submitted. In the future, when the dataset is large enough, we will split out a validation set that includes data from all colleges. The final learning rate we used is 5e-5. However, this can be changed according to different batch sizes when fine-tuning the model. The parameters we used are searched by a validation set from the engineering college documents (which takes the highest percentage of the total documents). The contrastive loss weight is 0.3, the alignment function weight is 0.1, and the number of candidates is 16. If people want to work in a different domain (other than College of Engineering), different parameters may have different performances. For the temperature in contrastive loss, we use the InfoNCE loss from van den Oord et al. (2018), and we use the default temperature value of 0.1. The temperature in decoding is 0.5. For model details shown in Figure 1, we use the same T5-encoder that shares parameters to keep our model DSPT5 as small (efficient) as possible.
@@
For ablation study setup, the model is trained on the entire training dataset without sampling. We only do the sampling on the evaluation dataset for the ablation study. Since the data split is 0.8:0.2 on each college domain, the evaluation set is also not a balanced dataset. In total, there are about 988 documents in the evaluation set, which is used for the evaluation in Table 2. After sampling, we have around 120 documents (which is now a balanced evaluation set covering all colleges) for the evaluation set ONLY used for the ablation study Table 3. We reported the average performance including all colleges in Table 3 for ablation study. Since the College of Engineering takes up almost half of the data shown in Table 1, the result of the average performance will be heavily influenced by the performance of the College of Engineering. Therefore, we sampled a balanced evaluation set to report the ablation study.
@@ Appendix D More Results
In this section, we provide the combined test set results of all eight colleges, which is shown in Tabel 9. From Table 9, we observe that our proposed DSPT5 outperforms all other baselines in the evaluation metric of s-BLEU, d-BLEU, BERTScore, BLONDE, ROUGE and METEOR. The closed-source model ChatGPT shows better performance in system-level evaluation metrics such as COMET. We conjecture that this is probably because such pre-trained LLMs (i.e. ChatGPT, Claude2, LLaMA2) are pre-trained on large quantities of documents, which make these models generate text more fluently, resulting in better system-level evaluation metrics. Alternatively, our proposed model DSPT5 is fine-tuned on our customized dataset VTechAGP, which results in better performance in word-level evaluation metrics.
@@ Appendix F Human Evaluation
To provide a comprehensive assessment of the generated general-audience abstract, we conducted a human evaluation involving our proposed model DSPT5 and all other baseline models using three independent experts  [^7] All judges have experience in scientific research and hold bachelor, master, and doctorate degree respectively.. Specifically, following a similar setting as Liu et al. (2024); Li et al. (2024); Song et al. (2024); Kew et al. (2023); Devaraj et al. (2021), our evaluation uses a random sample of 20 abstracts (10 in Computer Science and 10 in all other fields) from the test split VTechAGP considering the workload. Judges are presented with both the academic abstract and generated general-audience abstracts from seven models (DSPT5 and six baselines) for each data sample in total of 140 abstracts. Using a 1-5 Likert scale, the judges are asked to rate the model output based on five criteria: comprehensiveness, layness, meaning preservation, conciseness, and fluency. Details are shown in guidelines, which clarifies the meanings for each criteria in Figure 3 in the Appendix. Table 4 presents the average ratings from our human evaluation.
@@ Appendix G Methodology Explanation Details
In Sec. 4, we discuss the contrastive loss infoNCE derived from different learning representations. Here we provide a detailed explanation. To textual level, $r$ represents a set of keywords (here a set of keywords are textual keywords, and then $r$ is the representation (embedding) of the several textual keywords). As we said, $r^{pos}$ and $r^{neg}$ are not sets of vectors. Instead, $r^{pos}$ is a vector and $r^{neg}$ is a vector. They are the vector representation of a set of textual keywords. If there are $m$ keywords in the golden label and keyBERT returns $n$ keywords (where $m<n$), we sort the keywords by confidence and select the top $m$ keywords. Then we concatenate the words for dot production. For example if there are two terms in the label, we select two positive keywords from KeyBERT. The selection is based on the ranking of the confidence value provided by KeyBERT. Although dot production is sensitive to the order of the keywords, currently we do a simple concatenation and we leave the works for exploring new ranking approaches to make the keyword extraction module extract closer or more relevant words to the golden label in the future as the main contributions for this paper is about the new dataset VTechAGP and implementations for benchmark models.

### v2

@@ 6 Conclusion
We created VTechAGP  [^8] Benchmarks: https://github.com/SIGSEGV-0x7/VTechAGP-Benchmark, an academic-to-general-audience document-level translation benchmark dataset, which consists of academic and general audience abstract pairs with their corresponding auxiliary information such as title, subject terms, etc. We explore several SOTA LLMs to establish the baseline performance on the text paraphrasing task. We propose a new model, DSPT5, which includes dynamic soft prompt generation with hybrid loss during the training phase and a new crowd-sampling strategy in the inference stage. We evaluate the datasets and models from the perspective of document-level embedding-based, word-based, end-to-end metrics, simplicity, diversity, readability, toxicity, and human evaluation. Extensive experimental results on 8 different colleges of VTechAGP show that DSPT5 achieves comparable results with other SOTA benchmark models.
@@ 7 Limitations
We identify the following limitations (remaining challenges) and future directions:
(1) Dataset: Although VTechAGP is a recent and distinctive dataset, it suffers in terms of size. This is due to the fact that the policy of including a general audience version of the abstract for all ETDs submitted to VTechWorks began to be implemented in 2016. However, this aspect of the limitation should be gradually overcome as VTechWorks continues to collect and maintain this collection of documents. With the increase in graduate student enrollment at Virginia Tech in recent years, it is safe to assume that VTechAGP will continue to grow in the coming years. In addition, VTechAGP consists of pairs of abstracts carefully written by students, encapsulating their work during their time in graduate school. With about 50% of the corpus composed of Ph.D. dissertations and the other half composed of M.S. theses, we believe that VTechAGP should achieve good quality due to the amount of time and effort put into the creation process. As the size of the dataset grows, VTechAGP will become more robust for text paraphrasing tasks in the near future.
(2) Model: While transformer-based LLMs retain their popularity for text generation tasks, our experiment results in Sec. 5 show that there is still a huge room for further improvement of LLMs in academic to general audience text paraphrasing tasks. With the rapid development of LLMs, more robust and efficient LLMs are emerging. For example, the recently released model Gemini Team et al. (2023) shows great potential and competitive results compared to ChatGPT and LLaMA2 in natural language understanding-related tasks. In addition, some recently released lightweight decoder-only models (e.g., Phi-2 Li et al. (2023)) also exhibit great potential for text generation tasks. As the VTechAGP dataset grows, it will be able to support fine-tuning for more powerful open-source LLMs.
@@
Acknowledgments
This project was made possible in part by the Institute of Museum and Library Services (LG-256638-OLS-24).
@@ Appendix A Dataset Comparison
Breath-first search (BFS) is a fundamental building block in many graph-based applications, but it is difficult to optimize for a field-programmable gate array (FPGA) due to its irregular memory-access patterns. Prior work, based on hardware description languages (HDLs) and high-level synthesis (HLS), addresses the memory-access bottleneck of BFS by using techniques such as data alignment and compute-unit replication on FPGAs. The efficacy of such optimizations depends on factors such as the sparsity of target graph datasets. Optimizations intended for sparse graphs may not work as effectively for dense graphs on an FPGA and vice versa. This thesis presents two sets of FPGA optimization strategies for BFS, one for near-hypersparse graphs and the other designed for sparse to moderately dense graphs. For near-hypersparse graphs, a queue-based kernel with maximal use of local memory on FPGA is implemented. For denser graphs, an array-based kernel with compute-unit replication is implemented. Across a diverse collection of graphs, our OpenCL optimization strategies for near-hypersparse graphs delivers a 5.7x to 22.3x speedup over a state-of-the-art OpenCL implementation, when evaluated on an Intel Stratix 10 FPGA. The optimization strategies for sparse to moderately dense graphs deliver 1.1x to 2.3x speedup over a state-of-the-art OpenCL implementation on the same FPGA. Finally, this work uses graph metrics such as average degree and Gini coefficient to observe the impact of graph properties on the performance of the proposed optimization strategies.
@@ Appendix B Dataset Analysis
The original dataset does not include the college information. We manually assign each text sample to the college subset based on the department information on each official college website from Virginia Tech. Such data preprocessing steps introduce a few duplicates in the data because: (1) Some departments are involved in more than one college. For example, according to Virginia Tech’s official website, the Department of Biology Systems Engineering appears both in the College of Engineering  [^9] https://eng.vt.edu/academics/departments.html and the College of Agriculture and Life Sciences  [^10] https://www.cals.vt.edu/departments-and-school.html. (2) According to VT each College’s website, some departments (i.e. Environmental Science and Engineering, Counselor Education, etc.) can not be assigned to a specific college. We use fuzzy matches to find similar but not exact matches to assign data to colleges. Note that such a processing step may introduce a small number of duplicate files.
After processing, VTechAGP contains a total of 4,938 document pairs (academic abstract and general audience abstract), where 52.8% of the documents are from Ph.D. dissertations and 47.2% of the documents are from Master’s theses. Table 1 shows the data statistics of VTechAGP across eight different colleges. VTechAGP is divided into eight colleges because data from different colleges exhibits distinct statistics. In addition, the dataset divided into colleges can also support potential cross-domain research tasks and provide convenience for researchers focusing on a specific domain in the future. The document pairs are categorized based on colleges since authors from the same college may share some common terminology and background knowledge. Therefore, these categories should exhibit similar statistics. Table 1 reports some basic dataset statistics, including the number of documents for each college, the average number of sentences for each document, and the average sentence length for the source and target texts, respectively. We also report textual and lexical diversity (MTLD), which reflects the average number of words in a row for which a certain type-token ratio is maintained McCarthy (2005). To better illustrate the difference in readability between the source and target texts, we report Readability Consensus, [^11] https://pypi.org/project/textstat/ which is a combined evaluation of Flesch Kincaid Grade, Flesch Reading Ease, SMOG Index, Coleman-Liau Index, Automated Readability Index, Dale-Chall Readability Score, Linsear Write Formula, and Gunning FOG Formula. From Table 1, we observe that the general audience abstracts have fewer sentences and shorter sentence lengths compared to the academic abstracts. Furthermore, academic abstracts have better lexical diversity than general audience abstracts, except for the College of Architecture, Arts and Design, and the College of Business. In general, academic abstracts have a higher level of comprehension difficulty compared to general audience abstracts, except the College of Engineering. Analysis for the College of Engineering, which represents almost half of the total dataset, is discussed below.
@@
Department Readability Consensus
Aerospace Eng. 14th-15th / 14th-15th
Biomedical Eng. 13th-14th / 10th-11th
Building Construction 15th-16th / 15th-16th
Chemical Eng. 16th-17th / 11th-12th
Civil Eng. 16th-17th / 15th-16th
Computer Sci. 14th-15th / 15th-16th
Electrical & Computer Eng. 15th-16th / 15th-16th
Eng. Education 17th-18th / 16th-17th
Environmental Sci. & Eng. 15th-16th / 14th-15th
Industrial & Systems Eng. 16th-17th / 15th-16th
Materials Sci. & Eng. 16th-17th / 11th-12th
Mechanical Eng. 15th-16th / 14th-15th
Mining Eng. 15th-16th / 15th-16th
The distinct category (College of Engineering) contains the largest number of document pairs, representing almost half of the total dataset, due to the large number of students from 13 different departments in the College of Engineering. To better understand the characteristics of this category, we divided the documents in the College of Engineering based on their corresponding departments. Details are given in Table 8. Thus, we expect document pairs within this college to be more diverse compared to others. From Table 8, we observe that documents from most departments in the College of Engineering have a similar readability consensus between academic abstracts and general audience abstracts. The only outlier is the Department of Computer Science, which shows that general audience abstracts are more difficult to understand than academic abstracts. On the other hand, the Department of Electrical and Computer Engineering alone represents 15.2% of the total dataset, while maintaining an above-average readability consensus for its general audience abstracts, thus contributing to the higher readability consensus for the College of Engineering. We conjecture as follows: 1) Various complex terminologies from some research fields are inevitably retained in general audience abstracts because they cannot be replaced by simple terms. 2) Concepts in certain areas of engineering research are difficult to explain or rewrite into short and concise sentences. 3) Some engineering students are not as good at simplification writing as students from other colleges.
@@ Appendix C Parameter Settings
For the hyperparameters and configuration of DSPT5, we implement DSPT5 in PyTorch and optimize it with the AdamW optimizer. For pre-trained LLMs, we retrieve the generated output through their API by giving the prompt "Generate another version of the provided document for the general audience". We use grid search to tune the hyperparameters. The learning rate is $\in\left\{5e-2,5e-3,5e-4,5e-5\right\}$. The contrastive loss weight $\lambda$ in Equ. 8 and the alignment function weight $\gamma$ in Equ. 10 are chosen from 0.1 to 0.9 with a step size of 0.2. The number of candidates $C_{n}$ in Sec. 4.3 is chosen from $\left\{4,8,16\right\}$. The batch size is 4. The length of the source text, fixed prompt template, and keywords are 512, 16, and 16, respectively. The fixed prompt template used for baselines is: ‘Generate another version of the provided document for general audiences.’. We choose the hyperparameters based on the validation set. Because the dataset is not a balanced dataset, where College of Engineering takes up most of the datasets. We split a validation set in the domain of College of Engineering documents. This is because as shown in Table 1, it is not realistic to have a validation set in the Business domain, which has only 63 documents in total. The size of the dataset will only grow as new master’s theses or PhD dissertations are submitted. In the future, when the dataset is large enough, we will split out a validation set that includes data from all colleges. The final learning rate we used is 5e-5. However, this can be changed according to different batch sizes when fine-tuning the model. The parameters we used are searched by a validation set from the engineering college documents (which takes the highest percentage of the total documents). The contrastive loss weight is 0.3, the alignment function weight is 0.1, and the number of candidates is 16. If people want to work in a different domain (other than the College of Engineering), different parameters may have different performances. For the temperature in contrastive loss, we use the InfoNCE loss from van den Oord et al. (2018), and we use the default temperature value of 0.1. The temperature in decoding is 0.5. For model details shown in Figure 1, we use the same T5-encoder that shares parameters to keep our model DSPT5 as small (efficient) as possible. Here are the exact versions for LLM used: ChatGPT: gpt-3.5-turbo-0613 Claude2: Claude2.1 LLaMA: LLaMA2-7b BART: bart-large-cnn FlanT5: flan-t5-base T5: t5-base.
@@
For the ablation study setup, the model is trained on the entire training dataset without sampling. We only do the sampling on the evaluation dataset for the ablation study. Since the data split is 0.8:0.2 on each college domain, the evaluation set is also not a balanced dataset. In total, there are about 988 documents in the evaluation set, which is used for the evaluation in Table 2. After sampling, we have around 120 documents (which is now a balanced evaluation set covering all colleges) for the evaluation set ONLY used for the ablation study Table 3. We reported the average performance including all colleges in Table 3 for the ablation study. Since the College of Engineering takes up almost half of the data shown in Table 1, the result of the average performance will be heavily influenced by the performance of the College of Engineering. Therefore, we sampled a balanced evaluation set to report the ablation study.
@@ Appendix D More Results
In this section, we provide the combined test set results of all eight colleges, which is shown in Tabel 9. From Table 9, we observe that our proposed DSPT5 outperforms all other baselines in the evaluation metric of s-BLEU, d-BLEU, BERTScore, BLONDE, ROUGE and METEOR. The closed-source model ChatGPT shows better performance in system-level evaluation metrics such as COMET. We conjecture that this is probably because such pre-trained LLMs (i.e. ChatGPT, Claude2, LLaMA2) are pre-trained on large quantities of documents, which makes these models generate text more fluently, resulting in better system-level evaluation metrics. Alternatively, our proposed model DSPT5 is fine-tuned on our customized dataset VTechAGP, which results in better performance in word-level evaluation metrics.
@@ Appendix F Human Evaluation
To provide a comprehensive assessment of the generated general-audience abstract, we conducted a human evaluation involving our proposed model DSPT5 and all other baseline models using three independent experts  [^12] All judges have experience in scientific research and hold bachelor, master, and doctorate degree respectively.. Specifically, following a similar setting as Liu et al. (2024); Li et al. (2024); Song et al. (2024); Kew et al. (2023); Devaraj et al. (2021), our evaluation uses a random sample of 20 abstracts (10 in Computer Science and 10 in all other fields) from the test split VTechAGP considering the workload. Judges are presented with both the academic abstract and generated general-audience abstracts from seven models (DSPT5 and six baselines) for each data sample in a total of 140 abstracts. Using a 1-5 Likert scale, the judges are asked to rate the model output based on five criteria: comprehensiveness, layness, meaning preservation, conciseness, and fluency. Details are shown in guidelines, which clarify the meanings for each criterion in Figure 3 in the Appendix. Table 4 presents the average ratings from our human evaluation.
@@ Appendix G Methodology Explanation Details
In Sec. 4, we discuss the contrastive loss infoNCE derived from different learning representations. Here we provide a detailed explanation. To the textual level, $r$ represents a set of keywords (here a set of keywords are textual keywords, and then $r$ is the representation (embedding) of the several textual keywords). As we said, $r^{pos}$ and $r^{neg}$ are not sets of vectors. Instead, $r^{pos}$ is a vector and $r^{neg}$ is a vector. They are the vector representation of a set of textual keywords. If there are $m$ keywords in the golden label and keyBERT returns $n$ keywords (where $m<n$), we sort the keywords by confidence and select the top $m$ keywords. Then we concatenate the words for dot production. For example, if there are two terms in the label, we select two positive keywords from KeyBERT. The selection is based on the ranking of the confidence value provided by KeyBERT. We use T5 encoder, the inputs_embeds are in [batch_size, sequence_length, hidden_size], we concatenate embeddings in the sequence_length dimension. Although dot production is sensitive to the order of the keywords, currently we do a simple concatenation and we leave the works for exploring new ranking approaches to make the keyword extraction module extract closer or more relevant words to the golden label in the future as the main contributions for this paper is about the new dataset VTechAGP and implementations for benchmark models.